OpenStudy (anonymous):

if i subtract x^2 from, x^2 do i get x? or just no variable?

myininaya (myininaya):

anything-itself=0

OpenStudy (anonymous):

If you subtract a from a you get?

OpenStudy (anonymous):

u get zero

OpenStudy (anonymous):

No matter what a is (a=x^2 for example)

OpenStudy (anonymous):

okay, so i dont use the variable?

OpenStudy (anonymous):

And anyway, remember what I told you before. If you are adding or subtracting the exponents will not change.

myininaya (myininaya):

1x^2-1x^2=0*x^2

myininaya (myininaya):

but 0 times anything is 0

OpenStudy (anonymous):

but if you are dividing, you subtract exponents...thats what my problem is that i am working on

myininaya (myininaya):

1*x^2-1*x^2=1*(x^2-x^2)=1*0=0 does this look better to you

OpenStudy (anonymous):

\[x^2 - x^2\]\[=1x^2 - 1x^2\]\[=x^2( 1-1)\]\[=0x^2\]\[=0\]

myininaya (myininaya):

so pretty polpak

myininaya (myininaya):

i should had factor the the x^2 out instead darn it

OpenStudy (anonymous):

Please give the problem so we can understand what you mean lish. Are you subtracting x^2 from x^2 or are you dividing x^2 by x^2 ?

myininaya (myininaya):

\[\frac{x^2}{x^2}=1\]

myininaya (myininaya):

since \[1*x^2=x^2\]

OpenStudy (anonymous):

\[\frac{a}{a} = 1,\ \forall a \ne 0\]

myininaya (myininaya):

\[\frac{x^2}{x^2}=x^{2-2}=x^0=1\]

OpenStudy (anonymous):

sorry: the problem is 35x^5-35x^4+21x^2 / -7x^2

myininaya (myininaya):

\[\frac{35x^5-35x^4+21x^2}{-7x^2}?\]

OpenStudy (anonymous):

hey polpak after ur done helping this student please check out my question, thank you :)

OpenStudy (anonymous):

yes, thank you

myininaya (myininaya):

well i think you meant to write what i write i can read the way you right it but the i translated what you wrote is not what it should translate to what you wrote should translate to \[-35x^4+\frac{21x^2}{-7x^2}\]

myininaya (myininaya):

i suck at making sentences lol

myininaya (myininaya):

which is the correct translation?

OpenStudy (anonymous):

\[35x ^{5}-35x^4+21x^2 \div -7x^2\]

OpenStudy (anonymous):

\[-6x^{2}+35x^{5}-35x^{4}\]

OpenStudy (anonymous):

that is what i got